# Thermodynamic properties of the macroscopically degenerate tetramer-dimer phase of the spin-1/2 Heisenberg model on the diamond-decorated square lattice

### v1

@@ IV.3 Enhanced magnetocaloric effect due to the macroscopically degenerate DT phase
Finally, we discuss the magnetocaloric effect with a focus on the transition region between the macroscopically degenerate DT phase and the MD state (that is non-degenerate in a finite field). Figure 14 shows density plots of the entropy in the magnetic field versus temperature plane for the interaction ratio $J_{2}/J_{1}=1.7$. The superimposed constant-entropy curves correspond to the variation of the temperature $T$ during an adiabatic (de)magnetization process.
@@
For comparison, Fig. 14(b) presents the density plot of the entropy, for the same set of parameters and the same system size $N_{s}=40$, as obtained from Eq. (24) within the spin-star decoupling approximation. Remarkably, the spin-star approach is capable of predicting the aforementioned asymmetry of the entropy, and the relevant density plot reveals qualitatively the same features of the magnetocaloric effect as obtained from ED. Moreover, the enhanced magnetocaloric effect and cooling to absolute zero temperature should persist even in the thermodynamic limit due to the macroscopic degeneracy of the DT phase. The EMD model shows indeed that the asymmetric behavior of the isentropes as a function of magnetic field and temperature is preserved for a system size comprising of 16 elementary units in one spatial direction and infinite in the other direction, see Fig. 14(c). For this system size, the residual entropy of the DT phase is $s\approx 0.0587$, which is consistent with the macroscopic degeneracy of the hard-dimer model on the square lattice [64, 65, 43]. Hence, the enhanced magnetocaloric effect accompanied with cooling to zero temperature is detected for $h/J_{1}<0.3$ whenever the entropy is tuned below its zero-point value $s\lesssim 0.0587$ at the beginning of an adiabatic demagnetization process (i.e., all curves lying below the cyan curve).
@@ V Conclusion
In this study, we have investigated the thermodynamic properties of the spin-1/2 Heisenberg model on the diamond-decorated square lattice, particularly focusing on the macroscopically degenerate DT phase. Using a combination of exact diagonalization, sign-problem-free quantum Monte Carlo simulations, an effective monomer-dimer model, and a spin-star decoupling approximation, we have provided a comprehensive analysis of this model’s behavior under varying interaction ratios and external magnetic fields. Our results confirm that the DT phase exhibits unique thermodynamic properties due to its macroscopic degeneracy, which is linked to the classical hard-dimer model on the square lattice. The boundary between the MD and DT phases is well described by the effective monomer-dimer model, highlighting the significance of localized excitations in determining the low-temperature behavior of this system.
A particularly interesting result of our study is the enhanced magnetocaloric effect observed upon approaching the DT phase. This effect is particularly pronounced near the transition field between the MD and the DT phase, where the temperature decreases sharply during adiabatic demagnetization. This phenomenon arises due to the macroscopic degeneracy of the DT phase, which allows the system to cool to absolute zero temperature as long as the entropy content is set below the residual entropy value $s\approx 0.0583$ derived from the effective hard-dimer model on the square lattice in the thermodynamic limit. This renders the DT phase a promising candidate for achieving ultra-low temperatures by magnetic refrigeration. The third law of thermodynamics [78] requires this macroscopic degeneracy to be lifted in an experimental realization, for example by a lattice distortion that breaks the conservation of the total spin of a dimer. Our conclusions will nevertheless remain relevant as long as the resultant splitting of the ground-state manifold remains at an energy scale that is below the temperature scale of interest. The challenge is more to find a suitable compound to begin with. In this respect, it may be interesting to consider the diamond-decorated honeycomb lattice since in this case there are known compounds that exhibit the corresponding crystal structure, see for example Refs. [79, 80, 81].
Our findings demonstrate the usefulness of effective classical statistical physics models for the low-energy and low-temperature behavior of a highly frustrated quantum spin system and shed light on the rich phase diagram and thermodynamic phenomena of the spin-1/2 Heisenberg model on the diamond-decorated square lattice. Furthermore, they suggest potential applications of such highly frustrated quantum spin systems in magnetic cooling technologies.
@@ Acknowledgements.
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 945380. We acknowledge support under the Štefánik+ program for Slovak-France bilateral under contract SK-FR-22-0011/49880PG. Computing time for the QMC simulations was provided by the IT Center at RWTH Aachen University, part of the ED computations were carried out on the “osaka” cluster at the Centre de Calcul (CDC) of CY Cergy Paris Université. T.V. is supported through the EURIZON project (#3025 “Frustrated quantum spin models to explain the properties of magnets over wide temperature range”), which is funded by the European Union under grant agreement No. 871072. J.S. acknowledges financial support provided by Slovak Research and Development Agency under the contract No. APVV-20-0150. N.C. acknowledges support by the ANR through grant LODIS (ANR-21-CE30-0033).

### v2

@@ IV.3 Enhanced magnetocaloric effect due to the macroscopically degenerate DT phase
Finally, we discuss the magnetocaloric effect with a focus on the transition region between the macroscopically degenerate DT phase and the MD state (that is nondegenerate in a finite field). Figure 14 shows density plots of the entropy in the magnetic field versus temperature plane for the interaction ratio $J_{2}/J_{1}=1.7$. The superimposed constant-entropy curves correspond to the variation of the temperature $T$ during an adiabatic (de)magnetization process.
@@
For comparison, Fig. 14(b) presents the density plot of the entropy, for the same set of parameters and the same system size $N_{s}=40$, as obtained from Eq. (24) within the spin-star decoupling approximation. Remarkably, the spin-star approach is capable of predicting the aforementioned asymmetry of the entropy, and the relevant density plot reveals qualitatively the same features of the magnetocaloric effect as obtained from ED. Moreover, the enhanced magnetocaloric effect and cooling to absolute zero temperature should persist even in the thermodynamic limit because of the macroscopic degeneracy of the DT phase. The EMD model shows indeed that the asymmetric behavior of the isentropes as a function of magnetic field and temperature is preserved for a system size comprising of 16 elementary units in one spatial direction and infinite in the other direction, see Fig. 14(c). For this system size, the residual entropy of the DT phase is $s\approx 0.0587$, which is consistent with the macroscopic degeneracy of the hard-dimer model on the square lattice [64, 65, 43]. Hence, the enhanced magnetocaloric effect accompanied with cooling to zero temperature is detected for $h/J_{1}<0.3$ whenever the entropy is tuned below its zero-point value $s\lesssim 0.0587$ at the beginning of an adiabatic demagnetization process (i.e., all curves lying below the cyan curve).
@@ V Conclusion
In this study, we have investigated the thermodynamic properties of the spin-1/2 Heisenberg model on the diamond-decorated square lattice, particularly focusing on the macroscopically degenerate DT phase. Using a combination of exact diagonalization, sign-problem-free quantum Monte Carlo simulations, an effective monomer-dimer model, and a spin-star decoupling approximation, we have provided a comprehensive analysis of this model’s behavior under varying interaction ratios and external magnetic fields. Our results confirm that the DT phase exhibits unique thermodynamic properties owing to its macroscopic degeneracy, which is linked to the classical hard-dimer model on the square lattice. The boundary between the MD and DT phases is well described by the effective monomer-dimer model, highlighting the significance of localized excitations in determining the low-temperature behavior of this system.
A particularly interesting result of our study is the enhanced magnetocaloric effect observed upon approaching the DT phase. This effect is particularly pronounced near the transition field between the MD and the DT phase, where the temperature decreases sharply during adiabatic demagnetization. This phenomenon arises owing to the macroscopic degeneracy of the DT phase, which allows the system to cool to absolute zero temperature as long as the entropy content is set below the residual entropy value $s\approx 0.0583$ derived from the effective hard-dimer model on the square lattice in the thermodynamic limit. This renders the DT phase a promising candidate for achieving ultralow temperatures by magnetic refrigeration. The third law of thermodynamics [78] requires that this macroscopic degeneracy be lifted in an experimental realization, for instance through a lattice distortion that breaks the conservation of the total spin of a dimer. Our conclusions will nevertheless remain relevant as long as the resultant splitting of the ground-state manifold remains at an energy scale that is below the temperature scale of interest. The challenge is more to find a suitable compound to begin with. In this respect, it may be interesting to consider the diamond-decorated honeycomb lattice since in this case there are known compounds that exhibit the corresponding crystal structure, see for example Refs. [79, 80, 81].
While the present study does not directly address specific chemical realizations, our findings offer insights into the potential applications of frustrated spin systems, particularly in magnetic cooling technologies. These insights could be relevant for materials exhibiting similar thermodynamic properties, such as those found in the aforementioned compounds with a diamond-decorated honeycomb lattice.
Our findings demonstrate the usefulness of effective classical-statistical physics models for the low-energy and low-temperature behavior of a highly frustrated quantum spin system and shed light on the rich phase diagram and thermodynamic phenomena of the spin-1/2 Heisenberg model on the diamond-decorated square lattice. Furthermore, they suggest potential applications of such highly frustrated quantum spin systems in magnetic cooling technologies.
@@ Acknowledgements.
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 945380. We acknowledge support under the Štefánik+ program for Slovak-France bilateral collaboration under contract SK-FR-22-0011/49880PG. Computing time for the QMC simulations was provided by the IT Center at RWTH Aachen University, part of the ED computations were carried out on the “osaka” cluster at the Centre de Calcul (CDC) of CY Cergy Paris Université. T.V. is supported through the EURIZON project (No. 3025 “Frustrated quantum spin models to explain the properties of magnets over wide temperature range”), which is funded by the European Union under grant agreement No. 871072. J.S. acknowledges financial support provided by Slovak Research and Development Agency under the Contract No. APVV-20-0150. N.C. acknowledges support by the ANR through Grant LODIS (ANR-21-CE30-0033).